# Supernovae in 2023 (review): breakthroughs by late observations

### v3

@@ 4.1 SN Ia scenarios
The excess emission at early SN Ia phases (within a few days) is still puzzling (e.g., Lim et al. 2023; Ni et al. 2023a, b; Wang et al. 2023c). Processes might include the collision of the ejecta with a companion, with a close CSM, or the presence of ${}^{56}$i in the outer zones of the ejecta. Each of these has implications for the possible SN Ia scenarios. Wang et al. (2023c) conclude that no current model can adequately explain the full set of observations of the early emission excess of SN 2023bee. It seems that the early light excess cannot yet point to one leading SN Ia scenario, but has the potential to do so with more observations and their analysis.
The observation of SNe Ia during the nebular phase (e.g., Blondin et al. 2023; Camacho-Neves et al. 2023; Chen et al. 2023a; Deckers et al. 2023; Kumar et al. 2023; Kwok et al. 2023a; Liu et al. 2023d; Ni et al. 2023a) is critical to reveal the explosion mechanism as it can, for example, reveal the total mass of some isotopes that are nucleosynthesis at the explosion and their velocity distribution. Chen et al. (2023a) and Kwok et al. (2023a) deduce that SN 2021aefx has a spherically symmetric distribution of observed isotopes at the nebula phase. This supports the notion that many normal SNe Ia have large-scale spherical explosions (for review see Soker 2019). Blondin et al. 2023 argue that no existing SN Ia explosion model can fit all observed properties of SM 2021aefx.
@@
In analyzing observations one must bear in mind all SN Ia scenarios. This is evident in the study of SN 2020eyj. Kool et al. (2023) observed SN 2020eyj that they found to be an SN Ia that interacts with a helium-rich CSM. They attribute SN 2020eyj to the SD scenario with a helium non-degenerate donor star. Kool et al. (2023) specifically write that their observation “leave the SD scenario as the only viable alternative”. However, Soker & Bear (2023) propose a new channel of the CD scenario where there are two major CEE phases. In the second CEE phase, the helium-rich core expands to interact with a WD companion, forming a helium-rich CSM before the explosion. The explosion takes place with a MED time of weeks to tens of years.
@@ 4.2 The group of lonely white dwarf scenarios
One of my claims in this review is that many (but not all; see below) of the peculiar SNe Ia result from channels that have two stars at the explosion $N_{\rm exp}=2$, as I present in Table 1, while normal SNe Ia result mainly from the lonely WD group (e.g., Braudo & Soker 2023), namely, $N_{\rm exp}=1$. I concentrate here on two SNRs that were analyzed in 2023, SN 1181 which is a peculiar SN Iax, and SNR G1.9+0.3, which is probably a normal SN Ia. I first discuss some non-lonely WD scenarios.
@@ 4.2.1 Peculiar SNe Ia from $N_{\rm exp}=2$ scenarios
I present some recent studies that further motivate me to suggest that most peculiar SNe Ia result from $N_{\rm exp}=2$ scenarios, i.e., scenarios that have two stars at the time of the explosion.
From their observation of the super-Chandrasekhar peculiar SN Ia (03fg-like) SN 2020hvf during the nebular phase Siebert et al. (2023a) conclude that the explosion is asymmetrical with two velocity components. They discuss the DDet scenario in a double-degenerate system. Namely, one WD exploded and one survived. This suggestion, if holds, shows that a scenario with two stars at explosion can lead to an aspherical peculiar SN Ia. That paper also shows the potential of observations during the nebular phase in revealing the exploding system.
Kwok et al. (2023b) demonstrate the importance of observations during the nebular phase in their study of the peculiar (03fg-like or super-Chandrasekhar) SN 2022pul. The presence of carbon-oxygen rich CSM (Siebert et al., 2023b), the composition of the ejecta, and the asymmetry of the ejecta led Kwok et al. (2023b) to conclude that the best model of this peculiar SN Ia is the violent merger channel of the DD scenario. This channel has two WDs at the explosion, $N_{\rm exp}=2$. Li et al. (2023c) suggest the violent merger channel of the DD scenario for the peculiar SN 2016ije. In a different study of ‘super-Chandrasekhar’ SNe Ia, Fitz Axen & Nugent (2023) find the DD scenario to better fit these peculiar SNe Ia than an explosion of a massive magnetized WD.
The study by Hoogendam et al. (2023) is very relevant to the proposition that peculiar SNe Ia have different progenitors than most normal SNe Ia. Hoogendam et al. (2023) find that 2002es-like and 2003fg-like peculiar SNe Ia have different pre-peak UV color types of evolution compared to normal SNe Ia. They further suggest that these peculiar SNe Ia have different progenitors, and propose that these peculiar SNe Ia may originate in low-metallicity DD or CD scenarios enshrouded by a carbon-rich circumstellar medium.
@@ 4.2.2 SN 1181 and its SNR Pa 30
Schaefer (2023) presents a detailed discussion of SN 1181 and its SNR Pa 30. He proposes that the best scenario to match observations is a merger of oxygen-neon (ONe) and carbon-oxygen (CO) WDs (see also Ko et al. 2023), like in the simulations by Kashyap et al. (2018). However, the simulations by Kashyap et al. (2018) show that the outcome is a highly non-spherical ejecta, contrary to the structure of Pa 30. The requirement from scenarios of normal SNe Ia to be able to form spherical SNRs was the highlight of an earlier review of SNe Ia (Soker, 2019). This requires the system to relax after a binary interaction, whether a full merger or a mass transfer process. Namely, there must be a merger to explosion delay (MED) time. Maguire et al. (2023) study SN 2020udy which is another SN Iax and argue that it was an incomplete explosion of a near-$M_{\rm Ch}$ WD. For SN 2020udy, I would also claim for a MED time. The importance of MED time in normal SN Ia scenarios is the highlight of another earlier review (Soker, 2018).
Bogomazov & Tutukov (2023) study the DD scenario with CO WD and one ONe WD for normal SNe Ia (for different outcomes of the merger of a CO WD with a ONe WD, including a possible electron capture CCSN, see Wu et al. 2023). However, SN 1181 is a peculiar SN Ia. This and its spherical ejecta suggest a lonely WD at the explosion time, which resulted from either the CD scenario where either the core of the asymptotic giant branch star or the WD companion, which merged with the core, being ONe rich (a scenario that was studied by Canals, Torres, & Soker 2018). Another possibility is the DD-MED scenario where a CO WD and an ONe WD merged to form a near-$M_{\rm Ch}$ WD that exploded after it relaxed, hence the spherical explosion.
I conclude that SN 1181 with its large-scale spherical SNR suggests that it is a result of one of the lonely WD scenarios (Table 1).
@@
In Soker (2024a) I identified a point-symmetry in the morphology of SNR G1.9+0.3, the youngest known SNR in the Galaxy, which exploded around 1890-1900. I analyzed the X-ray image of this SNR that Enokiya et al. (2023) released and pointed at ten pairs of opposite clumps or filaments in the X-ray morphology, in addition to the opposite pair of ears that was noticed in the past. I present the analysis in Figure 5 with more details in the caption. Known SN Ia explosion models do not form such point-symmetric structures. I therefore concluded that the point-symmetric structure comes from the CSM which was a planetary nebula. Namely, SNR G1.9+0.3 is an SNIP, i.e., an SN inside a planetary nebula. In Soker (2024c) I discussed SNR G1.9+0.3 together with the brightest planetary nebulae in the Galaxy and proposed that the planetary nebula into which the progenitor of SNR G1.9+0.3 exploded had a multi-polar morphology.
Figure 5: The point-symmetric morphology of type Ia SNR G1.9+0.3. The original X-ray image with contours of CO emission and the notations of the ellipse and coordinates are from Enokiya et al. (2023). The double-headed arrows (DHA; red, white, and yellow), the notation of the ears, and the inset on the right panels with the notations of rotation and bending are from Soker (2024a). The left panel presents the basic point-symmetric morphology. Each of the lines DHA-a to DHA-f points at a pair of clumps (twin clumps). The group of twin pairs and the pair of ears form the point-symmetric morphology. Line DHA-$\tau$ is a tentative pair because the clumps are small and faint. The pairs of clumps that two white double-headed arrows point at are treated differently on the right panel. The right panel focuses on the two opposite ears and the arc at the base of each ear. The closed dashed-white line on the upper arc marks the boundary of the arc and the dashed black lines mark the three emission peaks within the arc. The boundary of the arc with its three peaks is rotated around itself by $180^{\circ}$ and matched to the arc at the base of the bottom (eastern) ear by a $9^{\circ}$ displacement as the inset indicates. Namely, there is a $9^{\circ}$ bend point symmetry of the two ears (DHA-0) and of the arc at the base of each ear (DHA-1 to DHA-3).
@@
The conclusion of the study of SNR G1.9+0.3 is that it was an explosion inside a planetary nebula of a near-$M_{\rm Ch}$ lonely WD, most likely in the frame of the CD scenario, with a tiny probability for the DD scenario.
@@ 4.2.4 Lonely WDs at the explosion time
For its novelty, the group of lonely WD scenarios deserves this subsection. As I indicate in Table 1 the group of lonely WD scenarios includes the DD-MED scenario and the CD scenario. The exact definition of a lonely WD explosion is one where at the time of explosion there is only one star in the system and it is dynamically relaxed. This implies that the explosion might be spherical, or axi-symmetrical if the lonely WD is rapidly rotating. No other material explodes but the WD itself, nor a companion nor an accretion disk.
@@
Observed properties are more complicated and not all explosions are spherical. For example, Ni et al. (2023a) deduce that SN Ia 2021aefx is an $M_{\rm Ch}$ explosion but the explosion is not spherical. In the lonely-WD scenarios, this is accounted for by an explosion of a rapidly rotating WD or by an off-center deflagration to detonation transition. The rotating WD is likely to sustain a strong magnetic field that might account for the high polarization of some emission lines of the descendant SN Ia. Yang et al. (2022) observed high polarization in the Ca II near-infrared triplet of the normal SN Ia 2021rhu, and propose that magnetic field is the cause of this high polarization.
@@
Consider first the DD-MED scenario. The dynamical relaxation timescale of a WD remnant of a merger is about several seconds. An accretion disk around the WD should also be depleted, which can take several minutes. Therefore, the MED time should be several minutes or longer. The accretion disk can blow a wind or launch jets forming a disk-originated matter (DOM). If the MED time is shorter than about a few days, the ejecta collides with the DOM within hours to a few days, leading to an early excess emission (e.g., Srivastav et al. 2023a for SN 2022ywc). Although the DOM might have a bipolar structure, the explosion can be spherical. If the MED time in the DD-MED scenario is much longer, then we can have a spherical explosion of a normal SN Ia, or a spherical mass ejection in a peculiar SN Ia that leaves a WD remnant after the explosion. The latter channel is a possible scenario for SN 1181 that formed the spherical SNR Pa 30 (Section 4.2.2).
@@
In the CD scenario, the MED time is longer than several days because the hydrogen-rich envelope that was ejected during the common envelope evolution must be at large distances for the explosion to be classified as SN Ia. In rare cases, the last envelope to be ejected is helium-rich (Soker & Bear, 2023). The CD scenario can explain SNe Ia that have massive CSM, up to $M_{\rm CSM}\approx 3M_{\odot}$. Since this CSM is or was ionized by the WD merger remnant, it is or was a planetary nebula at the time of the explosion, i.e., the SN Ia is an SNIP. As with many other planetary nebulae that are formed by binary interaction, the CSM might be elliptical or bipolar and possess a point-symmetric morphology. I suggest an SNIP in the CD scenario for SNR G1.9+0.3 (Section 4.2.3). The CD scenario with a MED time of $>10^{6}{~{}\rm yr}$, such that the CSM does not exist anymore, might also account for SN 1181 and its Pa 30 remnant.
@@ 5.1 Point symmetric supernova remnants (SNRs)
The research of SNRs is a major topic with tens of papers in 2023 (Section 2) that cover many aspects of SNRs. I concentrated mainly on SNR morphologies, and in particular on point-symmetric morphologies. As I described in Section 2, the point-symmetric morphology requires that either the explosion itself, as I argue for CCSNe, and/or the CSM, as I argue for SNe Ia, have point-symmetric morphology to begin with. The interaction of a non-point symmetric explosion with a non-point-symmetric CSM or ISM, cannot form point-symmetrical SNRs.

### v4

@@ 4.1 SN Ia scenarios
The excess emission at early SN Ia phases (within a few days) is still puzzling (e.g., Lim et al. 2023; Ni et al. 2023a, b; Wang et al. 2024b). Processes might include the collision of the ejecta with a companion, with a close CSM, or the presence of 56Ni in the outer zones of the ejecta. Each of these has implications for the possible SN Ia scenarios. Wang et al. (2024b) conclude that no current model can adequately explain the full set of observations of the early emission excess of SN 2023bee. It seems that the early light excess cannot yet point to one leading SN Ia scenario but has the potential to do so with more observations and analysis.
The observation of SNe Ia during the nebular phase (e.g., Blondin et al. 2023; Camacho-Neves et al. 2023; Chen et al. 2023; Deckers et al. 2023; Kumar et al. 2023; Kwok et al. 2023a; Liu et al. 2023d; Ni et al. 2023a) is critical to reveal the explosion mechanism as it can, for example, reveal the total mass of some isotopes that are nucleosynthesis at the explosion and their velocity distribution. Chen et al. (2023) and Kwok et al. (2023a) deduce that SN 2021aefx has a spherically symmetric distribution of observed isotopes at the nebula phase. This supports the notion that many normal SNe Ia have large-scale spherical explosions (for a review, see Soker 2019). Blondin et al. 2023 argue that no existing SN Ia explosion model can fit all observed properties of SM 2021aefx.
@@
In analyzing observations, one must bear in mind all SN Ia scenarios. This is evident in the study of SN 2020eyj. Kool et al. (2023) observed SN 2020eyj that they found to be an SN Ia that interacts with a helium-rich CSM. They attribute SN 2020eyj to the SD scenario with a helium non-degenerate donor star. Kool et al. (2023) specifically write that their observation “leave the SD scenario as the only viable alternative”. However, Soker & Bear (2023) propose a new channel of the CD scenario where there are two major CEE phases. In the second CEE phase, the helium-rich core expands to interact with a WD companion, forming a helium-rich CSM before the explosion. The explosion takes place with a MED time of weeks to tens of years.
@@ 4.2 The group of lonely white dwarf scenarios
One of my claims in this review is that many (but not all; see below) of the peculiar SNe Ia result from channels that have two stars at the explosion $N_{\rm exp}=2$, as I present in Table 1, while normal SNe Ia result mainly from the lonely WD group (e.g., Braudo & Soker 2024), namely, $N_{\rm exp}=1$. I concentrate here on two SNRs that were analyzed in 2023: SN 1181, which is a peculiar SN Iax, and SNR G1.9+0.3, which is probably a normal SN Ia. I first discuss some non-lonely WD scenarios.
@@ 4.2.1 Peculiar SNe Ia from $N_{\rm exp}=2$ scenarios
I present some recent studies that further motivate me to suggest that most peculiar SNe Ia result from $N_{\rm exp}=2$ scenarios, i.e., scenarios with two stars at the time of the explosion.
From their observation of the super-Chandrasekhar peculiar SN Ia (03fg-like) SN 2020hvf during the nebular phase Siebert et al. (2023) conclude that the explosion is asymmetrical with two velocity components. They discuss the DDet scenario in a double-degenerate system. Namely, one WD exploded, and one survived. This suggestion, if holds, shows that a scenario with two stars at explosion can lead to an aspherical peculiar SN Ia. That paper also shows the potential of observations during the nebular phase in revealing the exploding system.
Kwok et al. (2023b) demonstrate the importance of observations during the nebular phase in their study of the peculiar (03fg-like or super-Chandrasekhar) SN 2022pul. The presence of carbon-oxygen rich CSM (Siebert et al., 2024), the composition of the ejecta, and the asymmetry of the ejecta led Kwok et al. (2023b) to conclude that the best model of this peculiar SN Ia is the violent merger channel of the DD scenario. This channel has two WDs at the explosion, $N_{\rm exp}=2$. Li et al. (2023c) suggest the violent merger channel of the DD scenario for the peculiar SN 2016ije. In a different study of ‘super-Chandrasekhar’ SNe Ia, Fitz Axen & Nugent (2023) find the DD scenario to better fit these peculiar SNe Ia than an explosion of a massive magnetized WD.
The study by Hoogendam et al. (2023) is very relevant to the proposition that peculiar SNe Ia have different progenitors than most normal SNe Ia. Hoogendam et al. (2023) find that 2002es-like and 2003fg-like peculiar SNe Ia have different pre-peak UV color types of evolution compared to normal SNe Ia. They further suggest that these peculiar SNe Ia have different progenitors and propose that these peculiar SNe Ia may originate in low-metallicity DD or CD scenarios enshrouded by a carbon-rich circumstellar medium.
@@ 4.2.2 SN 1181 and its SNR Pa 30
Schaefer (2023) presents a detailed discussion of SN 1181 and its SNR Pa 30. He proposes that the best scenario to match observations is a merger of oxygen-neon (ONe) and carbon-oxygen (CO) WDs (see also Ko et al. 2023), like in the simulations by Kashyap et al. (2018). However, the simulations by Kashyap et al. (2018) show that the outcome is a highly non-spherical ejecta, contrary to the structure of Pa 30. The requirement from scenarios of normal SNe Ia to be able to form spherical SNRs was the highlight of an earlier review of SNe Ia (Soker, 2019). This requires the system to relax after a binary interaction, whether a full merger or a mass transfer process. Namely, there must be a merger to explosion delay (MED) time. Maguire et al. (2023) study SN 2020udy, which is another SN Iax, and argue that it was an incomplete explosion of a near-$M_{\rm Ch}$ WD. For SN 2020udy, I would also claim for a MED time. The importance of MED time in normal SN Ia scenarios is the highlight of another earlier review (Soker, 2018).
Bogomazov & Tutukov (2023) study the DD scenario with CO WD and one ONe WD for normal SNe Ia (for different outcomes of the merger of a CO WD with an ONe WD, including a possible electron capture CCSN, see Wu et al. 2023). However, SN 1181 is a peculiar SN Ia. This and its spherical ejecta suggest a lonely WD at the explosion time, which resulted from either the CD scenario where either the core of the asymptotic giant branch star or the WD companion, which merged with the core, being ONe rich (a scenario that was studied by Canals, Torres, & Soker 2018). Another possibility is the DD-MED scenario where a CO WD and an ONe WD merged to form a near-$M_{\rm Ch}$ WD that exploded after it relaxed, hence the spherical explosion.
I conclude that SN 1181, with its large-scale spherical SNR, suggests that it results from one of the lonely WD scenarios (Table 1).
@@
In Soker (2024a), I identified a point-symmetry in the morphology of SNR G1.9+0.3, the youngest SNR in the Galaxy, which exploded around 1890-1900. I analyzed the X-ray image of this SNR that Enokiya et al. (2023) released and pointed at ten pairs of opposite clumps or filaments in the X-ray morphology, in addition to the opposite pair of ears that was noticed in the past. I present the analysis in Figure 5 with more details in the caption. Known SN Ia explosion models do not form such point-symmetric structures. I therefore concluded that the point-symmetric structure comes from the CSM which was a planetary nebula. Namely, SNR G1.9+0.3 is an SNIP, i.e., an SN inside a planetary nebula. In Soker (2024c), I discussed SNR G1.9+0.3 together with the brightest planetary nebulae in the Galaxy and proposed that the planetary nebula into which the progenitor of SNR G1.9+0.3 exploded had a multi-polar morphology.
Refer to captionRefer to caption
Figure 5: The point-symmetric morphology of type Ia SNR G1.9+0.3. The original X-ray image with contours of CO emission and the notations of the ellipse and coordinates are from Enokiya et al. (2023). The double-headed arrows (DHA; red, white, and yellow), the notation of the ears, and the inset on the right panels with the notations of rotation and bending are from Soker (2024a). The left panel presents the basic point-symmetric morphology. Each of the lines DHA-a to DHA-f points at a pair of clumps (twin clumps). The group of twin pairs and the pair of ears form the point-symmetric morphology. Line DHA-$\tau$ is a tentative pair because the clumps are small and faint. The pairs of clumps that two white double-headed arrows point at are treated differently on the right panel. The right panel focuses on the two opposite ears and the arc at the base of each ear. The closed dashed-white line on the upper arc marks the boundary of the arc, and the dashed black lines mark the three emission peaks within the arc. The boundary of the arc with its three peaks is rotated around itself by $180^{\circ}$ and matched to the arc at the base of the bottom (eastern) ear by a $9^{\circ}$ displacement as the inset indicates. Namely, there is a $9^{\circ}$ bend point symmetry of the two ears (DHA-0) and of the arc at the base of each ear (DHA-1 to DHA-3).
@@
The conclusion of the study of SNR G1.9+0.3 is that it was an explosion inside a planetary nebula of a near-$M_{\rm Ch}$ lonely WD, most likely in the frame of the CD scenario, with a tiny probability for the DD-MED scenario.
@@ 4.2.4 Lonely WDs at the explosion time
For its novelty, the group of lonely WD scenarios deserves this subsection. As I indicate in Table 1 the group of lonely WD scenarios includes the DD-MED scenario and the CD scenario. The exact definition of a lonely WD explosion is one in which, at the time of the explosion, there is only one star in the system, and it is dynamically relaxed. This implies the explosion might be spherical or axisymmetrical if the lonely WD rapidly rotates. No other material explodes but the WD itself, nor a companion nor an accretion disk.
@@
Observed properties are more complicated, and not all explosions are spherical. For example, Ni et al. (2023a) deduce that SN Ia 2021aefx is an $M_{\rm Ch}$ explosion, but the explosion is not spherical. In the lonely-WD scenarios, this is accounted for by an explosion of a rapidly rotating WD or by an off-center deflagration to detonation transition. The rotating WD is likely to sustain a strong magnetic field that might account for the high polarization of some emission lines of the descendant SN Ia. Yang et al. (2022) observed high polarization in the Ca II near-infrared triplet of the normal SN Ia 2021rhu, and propose that magnetic field is the cause of this high polarization.
@@
Consider first the DD-MED scenario. The dynamical relaxation timescale of a WD remnant of a merger is about several seconds. An accretion disk around the WD should also be depleted, taking several minutes. Therefore, the MED time should be several minutes or longer. The accretion disk can blow a wind or launch jets, forming a disk-originated matter (DOM). If the MED time is shorter than about a few days, the ejecta collides with the DOM within hours to a few days, leading to an early excess emission (e.g., Srivastav et al. 2023a for SN 2022ywc). Although the DOM might have a bipolar structure, the explosion can be spherical. If the MED time in the DD-MED scenario is much longer, we can have a spherical explosion of a normal SN Ia or a spherical mass ejection in a peculiar SN Ia that leaves a WD remnant after the explosion. The latter channel is a possible scenario for SN 1181 that formed the spherical SNR Pa 30 (Section 4.2.2).
@@
In the CD scenario, the MED time is longer than several days because the hydrogen-rich envelope that was ejected during the common envelope evolution must be at large distances for the explosion to be classified as SN Ia. In rare cases, the last envelope to be ejected is helium-rich (Soker & Bear, 2023). The CD scenario can explain SNe Ia with massive CSM, up to $M_{\rm CSM}\approx 3M_{\odot}$. Since this CSM is or was ionized by the WD merger remnant, it is or was a planetary nebula at the time of the explosion, i.e., the SN Ia is an SNIP. As with many other planetary nebulae that are formed by binary interaction, the CSM might be elliptical or bipolar and possess a point-symmetric morphology. I suggest an SNIP in the CD scenario for SNR G1.9+0.3 (Section 4.2.3). The CD scenario with a MED time of $>10^{6}{~{}\rm yr}$, such that the CSM does not exist anymore, might also account for SN 1181 and its Pa 30 remnant.
The main theoretical challenge to the solid establishment of the group of lonely WD scenarios is to show that some core-WD and WD-WD merger remnants can live for up to $\approx 10^{4}{~{}\rm yr}$ and more before they explode (e.g., Soker 2022a). Very encouraging to this challenge is the study by Neopane et al. (2022), who argue that WD-WD mergers produce a substantial population within a narrow mass range close to the Chandrasekhar mass limit and that the MED time of these remnants might be as long as $\approx 100{~{}\rm yr}$ and more (see also Ilkov & Soker 2012). Neopane et al. (2022) study the process by which these near-$M_{\rm Ch}$-mass WD merger remnants lose angular momentum and undergo compression and spin-up. These might lead to an explosion. Future studies should further explore this process and explain longer MED times.
@@ 5.1 Point symmetric supernova remnants (SNRs)
The research of SNRs is a major topic, with tens of papers in 2023 (Section 2) that cover many aspects of SNRs. I concentrated mainly on SNR morphologies and, in particular, on point-symmetric morphologies. As I described in Section 2, the point-symmetric morphology requires that either the explosion itself, as I argue for CCSNe, and/or the CSM, as I argue for SNe Ia, have point-symmetric morphology to begin with. The interaction of a non-point symmetric explosion with a non-point-symmetric CSM or ISM cannot form point-symmetrical SNRs.